### v1

# Large Language Model-driven Meta-structure Discovery in Heterogeneous Information Network

Abstract.
Heterogeneous information networks (HIN) have gained increasing popularity in the recent decade for their ability to capture the complex relations between nodes of diverse types. Meta-structure was proposed to identify the important patterns of relations on HIN, which has been proven effective for extracting rich semantic information from HIN and facilitating graph neural networks to learn expressive representations. However, hand-crafted meta-structures pose challenges for scaling up to large and complex HINs, which draw wide research attention for developing automatic meta-structure search algorithms. Previous efforts predominantly concentrate on searching for meta-structures that have good empirical prediction performance, overlooking the explainability of discovered meta-structures. As a result, they often produce complex meta-structures that are prone to overfitting and incomprehensible to human researchers. To address this challenge, we draw inspiration from the emergent reasoning abilities of large language models (LLMs). In this paper, we propose a novel reasoning meta-structure search (ReStruct) framework that integrates LLM reasoning into the evolutionary procedure for meta-structure search. ReStruct uses a grammar translator to encode the meta-structures into natural language sentences, and leverages the reasoning power of LLMs to evaluate the semantically feasible meta-structures. Besides, ReStruct also employs performance-oriented evolutionary operations. These two competing forces allow ReStruct to jointly optimize for semantic explainability and empirical performance of meta-structures. Finally, we also design a differential LLM explainer that can produce natural language explanations for the discovered meta-structures, and refine the explanation by reasoning through the search history. Experiments on five representative HIN datasets demonstrate that ReStruct achieve state-of-the-art performance in both node classification and link recommendation tasks. Additionally, a survey study involving $73$ graduate students shows that the meta-structures and natural language explanations generated by ReStruct are substantially more comprehensible.
Heterogeneous Information Networks, Large Language Models, Graph Neural Networks.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Heterogeneous information networks (HINs) have been proven effective in jointly modeling network topology and multi-typed relations (Shi et al., 2016). As a result, HINs are increasingly adopted in a wide range of applications, including social media (Xu et al., 2019), information retrieval (Wang et al., 2019), and recommender systems (Han et al., 2020). To fully exploit the rich semantic information encoded in HINs, researchers proposed to use hand-crafted templates of relation sequences to model the complex proximities on HINs (Sun et al., 2011), which is known as meta-path. It was later extended to meta-structure to capture more general interaction patterns than linear paths. The recent advancement of heterogeneous GNNs showed that meta-structures can be leveraged to learn expressive representations for HINs (Xu et al., 2019; Wang et al., 2019), which are crucial for various downstream applications. However, the previous works on HINs modeling predominated leverage hand-crafted meta-structures, which heavily rely on the knowledge of domain experts. Therefore, they are difficult to scale up to larger and more complex HINs that are vital for real-world applications.
Motivated by the importance of domain adaptive meta-structure design, numerous recent research efforts are dedicated to designing algorithms for automatic meta-structure design. Researchers propose to use genetic algorithm (Han et al., 2020), deep reinforcement learning (Peng et al., 2021) and differentiable neural architectural search model (Ding et al., 2021) to automatically identify meta-structures that could lead to superior performance in heterogeneous GNNs. However, these previous attempts solely focus on the prediction performance of meta-structures, which often lead to highly complex structures that are difficult to explain and susceptible to overfitting. Such “meta-structures” deviate from the original inspiration of meta-structures research that aims to extract semantically clear features from HINs (Huang et al., 2016).
The recent breakthrough in large language models (LLMs) (Floridi and Chiriatti, 2020) offers a unique opportunity to address the challenges of meta-structure discovery. The scaled-up versions of LLMs show emergent abilities for a wide range of complex tasks that go beyond auto-regression token generation (Wei et al., 2022a). For example, researchers find chain-of-thoughts prompting can effectively unlock LLMs’ reasoning capability for commonsense, mathematical, and logical problems (Wei et al., 2022b). Such general-purpose reasoning capability holds huge potential to make sense of the rich semantic information in HINs, which could be path-breaking to current performance-oriented meta-structure search algorithms. That is, instead of searching for best-performing meta-structures, the reasoning power of LLM could facilitate the joint optimization for semantic explainability by emulating human reasoners. Such reasoning process could effectively prevent the overfitting of identified meta-structures, and produce human understandable knowledge from given HINs.
In this paper, we propose a novel ReStruct (short for Reasoning meta-Structure search) framework that integrates LLM reasoning into an evolutionary procedure for meta-structure search. Specifically, ReStruct employs a grammar translator to encode meta-structures with nested clauses in natural language (see Figure 2). It ensures the semantic meaning of meta-structures can be readily comprehended by LLMs as natural language sentences. Besides, we define a set of basic operations that can explore the feasible adjacent meta-structures of a given meta-structure, allowing ReStruct to search the adjacent possibilities of promising candidates. Different from pure performance-oriented search, ReStruct jointly evaluates the semantic feasibility and empirical performance to determine promising candidates. On one hand, ReStruct employs a few-shot LLM predictor to evaluate the semantic similarities between given meta-structure candidates and the searched meta-structures in past experience, and use a similarity-oriented LLM selector to identify the most promising candidates based on the semantic similarities. On the other hand, after the semantically feasible candidates are empirically evaluated with heterogeneous GNNs, ReStruct uses a evolutionary updater adopting the classic elimination-reproduction procedure to mimic natural selection processes that refine meta-structure candidates based on their performances. Finally, we design a differential LLM explainer that generates natural language explanations for the discovered meta-structure. It employs a chain-of-thought prompting technique to perform structural comprehension and performance attribution step-by-step. Such a reasoning process generates high-quality explanations by explicitly comparing the chosen meta-structures and the adjacent, not chosen ones.
We evaluate ReStruct on five representative HIN datasets. Experiments show that ReStruct achieve state-of-the-art performance on both node classification and link recommendation tasks. Besides, ReStruct can generate natural language explanations as it searches through the solution space. To effectively evaluate the explainability of discovered meta-structures, we perform a user survey on $73$ graduate students with domain knowledge of HIN research. Survey results show that the meta-structure discovered by our ReStruct method is chosen as the most comprehensible one by 46.6% participants against 3 strong baselines, outperforming the second best baseline by 61.8%. More importantly, the natural language explanations generated by our method are preferred by most human participants (77.6% on average) in a head-to-head comparison with baseline methods.
We summarize our main contributions below:
  • We propose a novel ReStruct framework that integrates LLM reasoning into an evolutionary meta-structure search procedure. ReStruct jointly optimizes the empirical prediction performance and semantic explainability of meta-structures, which coordinates the competing forces of empirical performance-oriented evolutionary updater and semantic similarity-oriented LLM selector. ReStruct represents a path-breaking development in meta-structure search algorithms, which could generate meta-structures that represent human digestible knowledge on HINs and are less prone to overfitting.
  • We design a grammar translator to encode meta-structures as natural language sentences, which unleashes the reasoning power of LLMs to make sense of the rich semantic information on HINs. On top of this, we design a differential LLM explainer that can generate human comprehensible natural language explanations for discovered meta-structures.
  • Extensive experiments show that ReStruct achieves state-of-the-art performance on five representative datasets. More importantly, a user survey validates that ReStruct substantially outperforms baseline methods in terms of the explainability of discovered meta-structures.
2. Preliminaries
Here, we provide the definitions of heterogeneous information networks, meta-paths, and meta-structures as in the literature.
Definition 3.1. Heterogeneous Information Network (HIN) (Sun et al., 2011). An information network (IN) is mathematically a graph denoted as $G=\{V,E,T,R,\sigma,\phi\}$, with $V=\{v_{1},v_{2},...,v_{n}\}$ being the set of nodes, $E=\{e_{1},e_{2},...,e_{m}\}$ being the set of edges, $T=\{t_{1},t_{2},...,t_{k}\}$ being the set of node types, and $R=\{r_{1},r_{2},...,r_{j}\}$ being the set of edge types. $\sigma:V\rightarrow T$ is a function that maps each node to its associated type, and $\phi:E\rightarrow R$ is a function that maps each edge to its associated type. The network schema of $G$ is then denoted as $S=\{T,R\}$. If $|T|>1$ (multiple types of nodes) or $|R|>1$ (multiple types of edges), $G$ is a heterogeneous information network (HIN). Otherwise, it is a homogeneous information network.
Definition 3.2. Meta-path (Sun et al., 2011). Given an HIN $G$, a meta-path $P=t_{1}\xrightarrow{e_{1}}t_{2}...\xrightarrow{e_{p-1}}t_{p}$, is a sequence of node types and edge types defined on the network schema $S$, connecting a single source node type and a single target node type. One meta-path may correspond to many meta-path instances in $G$.
Definition 3.3. Meta-structure (Huang et al., 2016). Given an HIN $G$, a meta-structure $T$ is a generalization of the meta-path to allow for the existence of graph structures between the source node type and the target node type.
3. Methods
In this section, we introduce our proposed methods in detail. In Section 3.1, we elaborate our novel design of natural language encoding of meta-structures to facilitate LLM’s understanding of its semantic meanings. In Section 3.2, we design three basic operations for generating candidate meta-structures. In Section 3.3, we design two LLM modules to evaluate and select candidate meta-structures with semantic similarity orientation. In Section 3.4, we combine LLM-guided optimization with evolutionary processes to form an effective derivative-free optimization framework. Figure 1 shows the overview of our framework.
Figure 1. Overview of our proposed ReStruct framework.
3.1. Natural Language Encoding of Meta-Structures
To facilitate the LLM’s comprehension, we design a grammar translator module to encode each meta-structure into a natural language sentence, as shown in Figure 2. For any meta-structure, we first traverse the structure to find all possible simple paths from the source node to the target node. Each resulting path is equivalent to a meta-path connecting the source and the target, decomposed from the original meta-structure. For each path, we encode it into a natural language sentence using nested clauses signified by THAT, which is a commonly-used grammar in English and thus hypothesized to be well-comprehended by the LLM. After obtaining the natural language encodings of the decomposed meta-path, i.e., sub-logics, we further combine them using a conjunction word AND to indicate the logical summation effect.
Figure 2. Natural language encoding of meta-structures.
3.2. Basic Operations for Candidate Meta-Structure Generation
To generate comprehensive candidates for LLM selection and meanwhile ensure validity, we define three basic operations for modifying any meta-structure, and design a set of components for these operations, analogous to playing Lego. Examples are shown in Figure 3.
  • INSERTION. This operation replaces one edge of the original meta-structure with a component.
  • GRAFTING. This operation takes a component, finds two nodes in the original meta-structure with the same type as the chosen component’s first and last node, and merges them respectively.
  • DELETION. This operation deletes a certain amount of nodes from the original meta-structure, and tries connecting all neighbors back.
In our experiments, we take all meta-paths with no more than $2$ nodes as components for INSERTION, and all meta-paths with no more than $3$ nodes as components for GRAFTING. DELETION does not require components as input, as it regards all existing nodes on the original meta-structure as operation candidates. Note that using components with more nodes can bring larger exploration space at each step, but may also result in chaos and confusion for the LLM, and thus we leave it as future work to investigate the optimal component settings.
Figure 3. Basic operations of exploring adjacent meta-structures.
3.3. Semantic Similarity-oriented LLM modules for Candidate Selection
To regulate the explainability of discovered meta-structures, we design two LLM modules that evaluate and select candidate meta-structures in a semantic similarity-oriented manner. We illustrate the interaction processes between the main program and the LLM agents in Figure 4.
3.3.1. Few-shot LLM Predictor
Following the execution of the aforementioned basic operations on each meta-structure, we derive a set of one-step neighbors as potential candidates. This collection comprises meta-structures that have been previously encountered and evaluated in earlier generations, as well as entirely novel ones. To inform the decision-making process with insights gleaned from previous evaluations, we design a few-shot LLM predictor to predict the performance $\hat{p}$ of each candidate based on a performance pool that stores the structures and corresponding performances of seen meta-structures. Additionally, the few-shot LLM predictor is asked to provide a self-estimated confidence value $\hat{c}$ for each prediction, resulting in a $(\hat{p},\hat{c})$ pair associated with each candidate. Intuitively, if the few-shot LLM predictor deems a candidate highly similar to a counterpart in the performance pool, it is inclined to assign higher confidence to this prediction. This is grounded in the notion that structural similarity is likely indicative of functional similarity. An illustrative prompt-response round is exemplified in Step 1 of Figure 4.
3.3.2. Similarity-oriented LLM Selector
Receiving a set of candidates and their corresponding predicted performances from the few-shot LLM predictor, we design an Similarity-oriented LLM selector to make the final decision of picking one single candidate to enter the next generation. During this process, the Similarity-oriented LLM selector is expected to jointly consider, and perhaps trade-off between, a meta-structure’s semantic meanings, the expected outcomes of choosing it, as well as the credibility of this outcome expectation. An illustrative prompt-response round is exemplified in Step 2 of Figure 4.
3.4. Performance-oriented Evolutionary Updater
With closed-source LLM modules in the loop, it is not feasible for us to obtain the gradient for optimizing meta-structure search. Therefore, we operationalize a derivative-free optimization framework with inspirations from the genetic algorithm. Specifically, we maintain a population of $N$ individuals, each representing a distinct meta-structure. In every generation, we first evaluate the performance of each meta-structure by using it to train a GNN for the given downstream task such as recommendation and node classification. After evaluation, the underperforming meta-structures are eliminated from the population. The surviving meta-structures undergo a reproduction phase, where duplication occurs with probabilities proportional to their performances. In essence, this phase uses promising meta-structures to replenish the population to its original size. Both elimination and reproduction processes mirror natural selection mechanisms that enable species evolution in the wild. After getting the modified population, we feed it into the LLM to improve each individual meta-structure. Under the evolutionary framework, this can be regarded as a way of targeted mutation, as in this process new nodes and/or edges can be generated and some of the existing nodes and/or edges will be removed. The modified population will be further evaluated at the onset of the next generation, forming a loop of derivative-free optimization.
3.5. Differential LLM Explainer
One prominent advantage of the LLM lies in its unparalleled ability for natural language generation. To harness this ability, we design a differential LLM explainer module to automatically generate human-comprehensible textual explanations elucidating the reasons behind the superior performance of discovered meta-structures. We design a prompting process in the chain-of-thought flavor (Wei et al., 2022b) to guide the differential LLM explainer in discerning the critical structural properties that contribute to performance enhancement. Specifically, the prompting process for analyzing a given meta-structure $T$ unfolds in the following two steps:
Step 1: Structural Comprehension. We first sample a set of $n$ one-step neighbors for $T$, and translate each of them into a natural language sentence according to Method 3.1. Then, we prompt the differential LLM explainer to conduct a comprehensive analysis of both the original meta-structure and all its sampled neighbors. This involves breaking down each of them into meaningful sub-structures and identifying the sub-structure functions.
Step 2: Performance Attribution. We first perform a quick evaluation of $T$ and each of the sampled neighbors separately by training a GNN with one structure at a time for downstream tasks. Then, we ask the differential LLM explainer to identify the presence/absence of beneficial/detrimental sub-structures in $T$. This involves a joint consideration of the evaluated performances and the structural analysis conducted in the previous step.
Combination of these two steps empowers the differential LLM explainer to unravel the intricate connections between structural properties and performance, providing a comprehensive understanding of the discovered meta-structures. The effectiveness of this module is further justified by a pilot study involving human evaluators, which is elaborated in Section 4.4.2.
Figure 4. Example of LLM prompts and feedbacks.
4. Experiments
4.1. Experimental Settings
4.1.1. Datasets
We evaluate our proposed framework ReStruct on two important tasks in heterogeneous information network learning: recommendation and node classification.
For the recommendation task, the goal is to predict the existence of links between source nodes (e.g., users) and target nodes (e.g., businesses or movies). Under this task, we conduct experiments on three widely used real-world datasets: Amazon, Yelp, and Douban  [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding. Each dataset includes user ratings of items, which are converted to 0-1 binary labels with a threshold of $2$ to indicate the presence or absence of preference. Among the user-item pairs with the label ’1’, we randomly select half of them as positive pairs, which are further randomly split into train-validation-test sets with a ratio of 3:1:1. The other half is reserved for network construction so as to prevent label leakage. We take all user-item pairs with the label ’0’ as negative pairs, and also randomly split them into train-validation-test sets to pair each positive pair. If the total number of negative pairs is not enough, we randomly sample unconnected items until reaching the desired number. We adopt AUC (the area under the ROC curve) as the evaluation metric.
For the node classification task, the goal is to predict the correct labels of nodes of a given type, such as the genre of a movie. Under this task, we conduct experiments on two other widely adopted real-world datasets: ACM and IMDB. In these datasets, the classification targeted node type is papers (subjects) for ACM and movies (genres) for IMDB, respectively. We follow the data splits used in previous works (Ding et al., 2021; Li et al., 2023; Yun et al., 2019). We adopt the Macro-F1 score as the evaluation metric.
Details of all five datasets can be found in Appendix A.
4.1.2. Baselines
We compare ReStruct with a set of state-of-the-art baselines. These baselines can be classified into three categories:
  • Hand-crafted meta-paths: HAN (Wang et al., 2019), which is a heterogeneous GNN that learns graph representation with multiple hand-crafted meta-paths and fuses them with a multi-head attention mechanism.
  • Automatically-searched meta-paths: RMSHRec (Ning et al., 2022), which adopts a reinforcement learning framework.
  • Automatically-searched meta-structures: (1) GEMS (Han et al., 2020), which searches meta-structures for recommendation on HINs with a genetic algorithm; (2) DiffMG (Ding et al., 2021), which formalized the problem into a neural architecture search flavor and search for meta-structures in a differentiable manner; (3) PMMM (Li et al., 2023), which further generalizes with multi-graph search.
4.1.3. Hyperparameter Settings
For our model, we run the algorithm for $30$ generations with a population size of $5$ and an elimination rate of $0.2$. When modifying each meta-structure, we randomly sample a set of $20$ candidates if there are too many of them from the one-step neighbors. When predicting meta-structure performances with the LLM predictor, we randomly sample a set of $30$ evaluation records from the performance pool to enable a few-shot learning paradigm through prompting with examples. For the LLM, we use the GPT-4 model by calling the OpenAI API. For all baseline models, we follow the implementation released by the authors. We fix the hidden dimensions to $64$ for all evaluated models, and tune other hyperparameters including learning rate, weight decay, and dropout by referring to the performances on the validation set. To reduce the noise brought by randomness during program execution, for each combination of (model, task, dataset), we run experiments with $10$ different random seeds and report the average performance with standard deviation.
4.2. Comparison on Recommendation
In Table 1, we report the experimental results of our proposed framework on the recommendation task compared to baselines. First, we observe that models using meta-paths (HAN, RMSHRec) exhibit substantially poorer performances than those using meta-structures (all the rest), confirming that the stronger expression capability of meta-structures is desired for heterogeneous graph learning. Second, ReStruct consistently achieves the best performance across three datasets, showcasing the effectiveness of our framework in identifying meaningful and useful structures in various HINs. Specifically, the performance gain over GEMS confirms that LLM-guided “targeted mutation” converges to better solutions that pure random mutation in a classic genetic framework.
HAN RMSHRec GEMS DiffMG PMMM ReStruct
Amazon 63.53$\pm$0.24 61.80$\pm$0.15 63.76$\pm$0.47 73.27$\pm$0.20 73.78$\pm$0.04 75.27$\pm$0.19
Yelp 74.54$\pm$0.09 59.39$\pm$0.81 75.46$\pm$0.28 77.63$\pm$0.40 76.76$\pm$0.17 84.04$\pm$0.24
Douban 85.12$\pm$0.12 79.72$\pm$0.32 90.84$\pm$0.10 93.94$\pm$0.04 94.31$\pm$0.02 94.49$\pm$0.03
Table 1. AUC (%) of recommendation on three datasets.
Figure 5. Discovered meta-structures on Yelp and the corresponding generated natural language explanations.
4.3. Comparison on Node Classification
In Table 2, we report the experimental results of our proposed framework on the node classification task compared to baselines. First, HAN and RMSHRec do not achieve desirable performances. Specifically, HAN is limited by its heavy reliance on hand-crafted meta-structures, while RMSHRec suffers from high instability presumably due to the limitation accompanied by the reinforcement learning framework. Second, DiffMG and PMMM demonstrate improved performances, showcasing the advantage brought by meta-structures over meta-paths, as well as the NAS searching framework. Finally, the meta-structure discovered by ReStruct achieves the best performance on both ACM and IMDB datasets.
HAN RMSHRec DiffMG PMMM ReStruct
ACM 91.20$\pm$0.25 90.32$\pm$0.74 92.65$\pm$0.15 92.76$\pm$0.14 92.82$\pm$0.23
IMDB 55.09$\pm$0.67 56.72$\pm$2.60 61.04$\pm$0.56 61.69$\pm$0.40 63.32$\pm$0.62
Table 2. Macro F1 scores (%) of node classification on two datasets.
4.4. Explainability Analysis of Discovered Meta-Structures
4.4.1. Visualization of Discovered Structures and LLM-generated Explanation
In Figure 5, we present $3$ discovered meta-structures that are among the top-performing ones on the Yelp dataset for the recommendation task, along with a summary text explaining the structural attributes underlying their outstanding performances, which are essentially output from our differential LLM explainer further condensed for clarity. We highlight the LLM-identified good (relevant) sub-structures in red, while the bad (distracting) sub-structures in blue. This visualization facilitates a comprehensive understanding of each structure’s composition. For example, the meta-structure in Figure 5 (a) contains critical yet simple sub-structures describing the geographical, social, and business category contexts of user behavior. While more nodes and edges can create complex relationships, they are not always desirable for HIN learning, as showcased by the identified distracting sub-structures. These structures, though challenging to handcraft, carry semantically meaningful explanations that remain accessible with textual assistance.
4.4.2. Pilot Study of Human Evaluation
We conduct a pilot study with $73$ human participants, more specifically, graduate students with domain knowledge of HIN research, to evaluate the level of explainability provided by our framework compared to baselines. We structure the study around two sets of questions. The first set aims to compare the explainability of the best-performing meta-structure discovered by our model with those found by several baseline models, including GEMS, DiffMG, and PMMM. Participants are asked to select one meta-structure from a set of meta-structures, each being the best one found by a specific model, as documented in the original papers (Han et al., 2020; Ding et al., 2021; Li et al., 2023). The second set of questions investigates the comprehension gain brought by the textual explanation generated by our differential LLM explainer, as described in Method 3.5, coined the Differential Explanation. As a baseline, we include the textual explanation generated by directly prompting an LLM to explain the reasons behind a given meta-structure’s strong performance, without undergoing the two-step prompting process, coined the Non-Differential Explanation. For three meta-structures among the best ones discovered by our model on the Yelp HIN, participants are presented with both types of explanations. They are then asked to determine which explanation is more helpful in enhancing their (1) understanding of how this meta-structure is constructed, and (2) insights on how to design a better one. The complete questionnaire utilized in our study is available in Appendix B.
Figure 6 shows the result of our first question set. Among all four models, the meta-structure discovered by our model is regarded by most people ($46.6\%$) as the most informative or comprehensible one, outperforming the second baseline, DiffMG ($28.8\%$), by $61.8\%$. GEMS and PMMM follow with the same level of recognition, $12.3\%$. This outcome aligns with our initial objective of discovering meta-structures that not only excel in downstream tasks but are also accessible and easily understood by human users, presumably because our framework seeks to find meta-structures that are semantically meaningful while keeping as simple as possible.
Figure 7 shows the result of our second question set. When comparing two types of generated textual explanations, the Differential Explanation is consistently and significantly more preferred by human participants ($77.6\pm 2.8\%$) across various meta-structures. This outcome justifies the effective design of our differential LLM explainer that unleashes the LLM’s reasoning ability on the intricate connections between sub-structures, sub-functions, and how they interact to determine the ultimate performances.
Figure 6. Human evaluated explainability of the best meta-structures discovered by different models.
Figure 7. Explainability gain brought by LLM-generated differential meta-structure explanation compared to non-differential explanation.
4.5. Ablation Study
To assess the efficacy of the designed grammar translator and LLM decision modules, we conduct ablation studies involving two model variations: (1) No Grammar: In this scenario, the grammar translator module is excluded from the framework. LLM agents are prompted to predict candidate performances and select the best candidates directly based on the matrix expression of meta-structures (Han et al., 2020). (2) No Grammar, No LLM Decision: This variation involves the removal of both the grammar translator module and the LLM decision modules from the framework. Candidates are randomly selected to advance to the next generation.
As shown in Figure 8, both ablation models exhibit inferior performance compared to the full model. Specifically, among the three models, No Grammar, No LLM Decision is more prone to getting stuck in local sub-optimal solutions. The LLM selection module facilitates a faster escape from the sub-optimal, showcasing the advantage of knowledge-guided exploration. Ultimately, the full model navigates the search space with significant strides, suggesting that our grammar encoding of the meta-structures contributes to performance enhancement by enabling LLM agents to better grasp the semantic meanings of complex meta-structures. In summary, these results affirm that both designed components contribute considerably to a better meta-structure discovery process.
Figure 8. Comparison of best-performing meta-structure found by each generation.
5. Related Works
5.1. Identifying Meta-structure in HIN
In the past decade, heterogeneous information networks (HIN) gradually gained popularity for their ability to capture the complex relations between multi-typed nodes, which play important roles in various research areas such as information retrieval and social network modelling (Shi et al., 2016). Meta-path, a pre-defined path template of relation sequences, was proposed to measure the similarities between nodes on HIN (Sun et al., 2011). It allows search algorithms like PathSim (Sun et al., 2011) to find peer nodes that are connected by paths with different semantic meanings. The concept of meta-path was later extended to a more general form of meta-structure (Huang et al., 2016), where the relation patterns between connected nodes can be characterized as a directed acyclic graph. Previous works have found meta-structures useful for boosting machine learning performance on HINs (Jiang et al., 2017). However, early works were based on carefully hand-crafted meta-structures, which heavily relied on domain knowledge. To address this challenge, several more recent works proposed to automate meta-structure design with heuristic algorithms (Meng et al., 2015), reinforcement learning (Yang et al., 2019), evolutionary search (Han et al., 2020) and differentiable structure learning (Ding et al., 2021). Different from previous efforts of automatic meta-structure design, we are first to leverage the emergent reasoning ability of LLMs (Wei et al., 2022a). We design novel LLM modules for the automatic generation, evaluation, and explanation of novel meta-structures, which are proven effective in several representative datasets.
5.2. Deep Learning on HIN
The success of graph neural networks introduces revolutionary deep learning techniques into HIN modeling (Wu et al., 2020). Metapath2Vec (Dong et al., 2017) proposed to learn deep representations for nodes in HIN via meta-path-guided random walks. Attention mechanism was later introduced to learn more expressive representations for HIN, which were proven effective for link prediction (Xu et al., 2019) and node classification (Wang et al., 2019). Extensive follow-up research was devoted to designing more effective heterogeneous GNN frameworks (Li et al., 2021a; Fu et al., 2020). Besides, considerable research efforts were drawn to replace handcrafted meta-structures with automatic search. GEMS (Han et al., 2020) proposed to combine heterogeneous GNN with evolutionary algorithms to identify meta-structures and learn deep neural networks simultaneously. Several deep reinforcement learning models and neural architecture search models are also proposed to jointly optimize the meta-structures with heterogeneous GNN (Ding et al., 2021; Peng et al., 2021; Ning et al., 2022; Li et al., 2023; Wan et al., 2020; Li et al., 2021b). However, previous automatic meta-structure design method solely focused on prediction performance, often yielding complex and difficult to explain meta-structures. To the best of our knowledge, our study is the first to harness the semantic reasoning capability of LLMs for automatic meta-structure design. Our model can discover meta-structures that not only show high prediction performance, but also can be adequately explained with natural languages.
5.3. LLM for Graph Learning
The recent advancement of LLMs shows general capabilities that go beyond natural language tasks. Researchers on graph learning are particularly interested in their ability to graph reasoning tasks (Guo et al., 2023a). For example, there are several recent attempts to use LLMs to enhance node features or employ them as standalone graph predictors (Chen et al., 2023). The research communities are also actively discussing the perspective for developing large graph models (Zhang et al., 2023; Liu et al., 2023). Recent research found LLMs exhibited preliminary ability for graph learning tasks, such as detecting connectivity and cycles, performing topological sort, and emulating GNNs (Wang et al., 2023). Besides, LLMs can also effectively perform reasoning tasks on knowledge graph (Sun et al., 2023). To better align graph problems with LLM, several recent works propose various methods to encode the geometric structure and node features of graph problems (Zhao et al., 2023; Fatemi et al., 2023). On top of these textual encodings, researchers have explored the possibility of replacing GNNs with LLM reasoning (Ye et al., 2023) and performing instruction tuning for graph tasks (Tang et al., 2023). In this paper, we fundamentally extend LLM reasoning to HIN meta-structure discovery. Specifically, we propose a novel meta-structure encoding method, which effectively boosts LLM’s reasoning capability on HINs.
5.4. LLM for Pattern Discoveries
The scaled-up language models have emerged reasoning capability for general tasks (Wei et al., 2022b), including commonsense reasoning, logical reasoning, and mathematics reasoning. Optimized prompting routines like chain-of-thought (CoT) (Wei et al., 2022b) and tree-of-thoughts (ToT) (Yao et al., 2023) can effectively improve the scaling curve of LLM’s reasoning capability. As a result, recent research showed that LLM can be leveraged to identify novel patterns and feasible solutions in a large space. For example, FunSearch was proposed for mathematical algorithms discovery (Romera-Paredes et al., 2023). Besides, previous works also designed novel LLM-driven algorithms for evolutionary search (Guo et al., 2023b), reinforcement learning (Shinn et al., 2023), and hyper-parameter optimization (Liu et al., 2024). In this paper, we propose a novel framework to harness the reasoning power of LLM for meta-structure discovery. Our framework equips LLM with enhanced capability to understand the semantic meaning of meta-structures and search for promising candidates.
6. Conclusion
In this work, we propose a novel framework, ReStruct, that fuses the power of LLMs with evolutionary algorithms to facilitate automatic meta-structure discovery across diverse HINs. Extensive experiments demonstrate that our model excels in uncovering previously undiscovered meta-structures, thereby significantly enhancing downstream model performance compared to state-of-the-art baselines. Notably, a pilot study involving human participants confirms that our model provides a substantial level of interpretability regarding the identified meta-structures. For future work, we will explore the feasibility of finetuning local models to mitigate network communication costs associated with API calls.
Appendix A Statistics of Datasets
Figure 9. Statistics of datasets for node classification
Figure 10. Statistics of datasets for recommendation
Appendix B Questionnaire for Meta-Structure Explainability
Please find the questionnaire in the next few pages.